OpenStudy (anonymous):

write an equation in the form y=mx+b for the line which contains the point (6,-8) and (-2,-12)

OpenStudy (anonymous):

i need help

OpenStudy (mrnood):

do you know what the form of the equation y=mx+b means?

OpenStudy (anonymous):

no i do not

OpenStudy (mrnood):

so - you are studying straight line equations have you seen the equation y=mx + b before?

OpenStudy (anonymous):

yes i have but i honestly dont really get it

OpenStudy (mrnood):

do you know what the slope of a straight line is?

OpenStudy (anonymous):

are you talking bout the y1,y2/x1,x2

OpenStudy (mrnood):

yes - that is correct (but it is y1-y2/x1-x2) you can use the 2 points you are given to work out the slope from that equation

OpenStudy (mrnood):

fill in the numbers - from the points: x1= x2= x1-x2= y1= y2= y1-y2= y1-y2/x1-x2 =

OpenStudy (anonymous):

ok give me a min im going to do it

OpenStudy (anonymous):

my answer is 0.5

OpenStudy (mrnood):

that is good - correct now in the equation y=mx +b the slope is the value m so you can start to make the equation y= 0.5x +b OK so far?

OpenStudy (anonymous):

so should i make -12 or -8 my y intercept

OpenStudy (mrnood):

so - don't rush ahead... the equation is y=0.5x+b and you know that the point (6,-8) is on th eline so you can fill in some numbers: -8 = 0.5(6) +b (that is th esame equation but we pput the values from one of the known points in OK? so now you can work out b from that equation...

OpenStudy (mrnood):